The parallel plates of capacitors are 5 cm x 10 cm and have a separation of 2cm. If the space between them the plates is filled with bakalite (k=4.7) what is the capacitance?


1
Expert's answer
2021-04-19T17:09:56-0400

The capacitance of a parallel plate capacitor is given as follows:


"C = \dfrac{\varepsilon_0\varepsilon A}{d}"

where "\varepsilon_0 = 8.85\times 10^{-12}F/m" is the vacuum permittivity, "\varepsilon = 4.7" is the bakalite dielectric permittivity, "A = 5cm\times 10cm = 50cm^2 = 0.005m^2" is the area of one plate, and "d = 2cm = 0.02m" is the distance between the plates. Thus, obtain:


"C = \dfrac{8.85\times 10^{-12}F/m\cdot 4.7 \cdot 0.005m^2}{0.02m} \approx 1.04\times 10^{-12}F"

Answer. "1.04\times 10^{-12}F".